OpenStudy (anonymous):

in general how would you factor the polynomial (a^2-b^2)

OpenStudy (anonymous):

(6x+5)(6x-5)

OpenStudy (mathstudent55):

correct

OpenStudy (goformit100):

it is still incorrect

OpenStudy (goformit100):

Where has '2' gone ?

OpenStudy (mathstudent55):

the 2 is a sqaure

OpenStudy (anonymous):

Nice job @romanortiz65 ... you have been working hard at learning to factor.

OpenStudy (goformit100):



OpenStudy (goformit100):

hey 2 is not square

OpenStudy (goformit100):

2 (6)^2 - (5)^2

OpenStudy (anonymous):

dude its x^2

OpenStudy (goformit100):

i think it will be Ans : 2(6x+5)(6x-5)

OpenStudy (goformit100):

it is Question Error

OpenStudy (mathstudent55):

@goformit100 Your factoring is incorrect.

mathslover (mathslover):

\(36x^2 - 25 = (6x)^2 - 5^2 \) Using : \(a^2 - b^2 = (a+b)(a-b)\) So, \((6x+5)(6x-5) \)

OpenStudy (goformit100):

then you must write '36x^2-25' instead '36x2-25'

OpenStudy (mathmate):

36x2-25 is meant to be \( 36^2 -25 \).

OpenStudy (anonymous):

okay sleek-feathered one no ones perfect okay

mathslover (mathslover):

@goformit100 , but still if you say : \(2(6x+5)(6x-5) = 2(36x^2 - 25) = 72x^2 - 50\) So in any case, you have made mistake in factroing.

OpenStudy (goformit100):

ya if it x^2 then it is correct ..... Well done @mathslover

OpenStudy (mathmate):

Sorry, I meant \( 36x^2 -25 \)

OpenStudy (mathstudent55):

@goformit100 Where have you been for the last several years? You are correct, but haven't you seen countless people write 5x2 when they mean 5x^2?

OpenStudy (mathstudent55):

@goformit100 No matter what the OP meant, your factoring is still not correct.